OpenStudy (anonymous):

simplify sqrt 96/sqrt 8

OpenStudy (anonymous):

rationalize denominator = first step

OpenStudy (anonymous):

sqrt(96)=sqrt(16)sqrt(6)=4sqrt(6)=4sqrt(3)*sqrt(2) sqrt(8)=sqrt(4)*sqrt(2)=2sqrt(2) so: \[\frac{4\sqrt{3}\sqrt{2}}{2\sqrt{2}}\]

OpenStudy (anonymous):

you can simplify using division if they are both rationals of the same degree

OpenStudy (anonymous):

I get 2 times the square root of 3.

OpenStudy (anonymous):

thats correct by my math

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

so which answer is correct?

myininaya (myininaya):

that's right gj jah and evel

OpenStudy (anonymous):

glad to help.

OpenStudy (anonymous):

96 /8 =12 and simplify

OpenStudy (anonymous):

\[\frac{\sqrt{96}}{\sqrt{8}}=\sqrt{\frac{96}{8}}=\sqrt{12}=\sqrt{4}*\sqrt{3}=2\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{96*8}/8\]\[\sqrt{768}/8\]

OpenStudy (anonymous):

sorry really bad drawing

OpenStudy (anonymous):

its okay thanks